### v1

# ChemKANs for Combustion Chemistry Modeling and Acceleration

Abstract
Efficient chemical kinetic model inference and application for combustion problems is challenging due to large ODE systems and wideley separated time scales. Machine learning techniques have been proposed to streamline these models, though strong nonlinearity and numerical stiffness combined with noisy data sources makes their application challenging. The recently developed Kolmogorov-Arnold Networks (KANs) and KAN ordinary differential equations (KAN-ODEs) have been demonstrated as powerful tools for scientific applications thanks to their rapid neural scaling, improved interpretability, and smooth activation functions. Here, we develop ChemKANs by augmenting the KAN-ODE framework with physical knowledge of the flow of information through the relevant kinetic and thermodynamic laws, as well as an elemental conservation loss term. This novel framework encodes strong inductive bias that enables streamlined training and higher accuracy predictions, while facilitating parameter sparsity through full sharing of information across all inputs and outputs. In a model inference investigation, we find that ChemKANs exhibit no overfitting or model degradation when tasked with extracting predictive models from data that is both sparse and noisy, a task that a standard DeepONet struggles to accomplish. Next, we find that a remarkably parameter-lean ChemKAN (only 344 parameters) can accurately represent hydrogen combustion chemistry, providing a 2 acceleration over the detailed chemistry in a solver that is generalizable to larger-scale turbulent flow simulations. These demonstrations indicate potential for ChemKANs in combustion physics and chemical kinetics, and demonstrate the scalability of generic KAN-ODEs in significantly larger and more numerically challenging problems than previously studied.
keywords:
Kolmogorov-Arnold Networks, Chemical Kinetics, Dynamical Systems , Reacting Flows , Model Inference , Computational Acceleration1 Introduction
Chemical kinetic modeling is a key backbone of many areas of science and engineering, from biological systems to energy conversion. Traditional hurdles to effective kinetic modeling for combustion have included large and complex reaction mechanisms that can be challenging to derive manually via expert knowledge [1], and the substantial size and stiffness of these chemical kinetic systems [2, 3] leading to large computational cost when integrating their solutions. Recent developments in machine learning have shown promise in helping to address these two issues [4] through streamlined model inference techniques and efficient, data-driven solvers.
Machine learning tools have been proposed using diverse algorithms, presumed model forms, and optimization approaches to accomplish automatic kinetic model discovery from various types of experimental data. The Chemical Reaction Neural Network approach [5], for example, is capable of inferring reaction networks and parameters from limited species trajectory or heat release data [6, 7, 8] by directly enforcing the Arrhenius and mass action laws in a neural network structure. The Sparse Identification of Nonlinear Dynamics (SINDy) approach is similarly capable of extracting models from experimental data by assuming various functional relationship building blocks and learning the precise forms needed to fit the data [9]. Further optimization and inverse modeling tools exist for other chemical kinetic inference problems [10, 11, 12], with a key piece of many physics-based model inference techniques being a certain (and often substantial) degree of prior knowledge of the governing equations, reaction pathways, and reactants. A critical capability required of such kinetic model discovery algorithms, especially in the field of combustion, is robustness to noisy data or uncertain model pathways and parameters [5, 13, 7], with realistic models often containing significant uncertainty on the kinetic parameters.
On the solver side, dimension reduction [6, 14] and computational acceleration [15, 16, 17, 18] are two general tools that aim to address the issues of large mechanisms and slow stiff solvers. For instance, Owoyele and Pal [17] recently proposed ChemNODE, a creative and high-performing tool which leverages the neural ODE concept of Chen et al. [19] to replace a complete chemical kinetic model with a collection of neural networks, one for each tracked thermochemical quantity. By directly linking the current thermochemical state to the chemical source term with no other problem-specific treatment, computational acceleration was enabled in the as-studied homogeneous reactor while retaining the generalizability of the surrogate model to higher-dimensional reacting flows where such acceleration becomes significantly more meaningful. Other recent works leverage DeepONets [20] to directly learn stiff integrators using neural operators, either with problem-specific network structures [21, 14] or by mapping from the current state to the source term, like in ChemNODE [22, 18], allowing for significant computational acceleration downstream. These strengths all come with drawbacks, however. Owoyele and Pal [17], for example, found that while the neural ODE approach’s clever exploitation of the dynamical structure of chemical kinetic models can provide high accuracy, the nonlinearity inherent to such models creates a challenging inference problem for the underlying MLP layers that led the authors to omit a handful of species (including the key H radical) and break the training up into multiple unique and likely redundant networks, rather than a single cohesive architecture. DeepONet techniques are cheap to evaluate and capture steady state behavior well, but their accuracy can suffer in stiff regions of the data that the integrator (which with DeepONets must be inferred directly by the network, as they do not explicitly leverage existing ODE solvers as Neural ODE approaches do) can learn to skip over without significant penalties. We thus find that despite these recent novel and productive efforts, the training of efficient surrogate models for combustion chemistry remains a challenging task with open questions due to stiff behavior in the solution profiles and numerical instability in the nonlinear training processes.
Kolmogorov-Arnold Networks were proposed recently [23] as an alternative to multi-layer perceptrons for general neural network applications, where instead of learning weights and biases on fixed activation functions, the shapes and magnitudes of the activation functions themselves are learned via gridded basis function sums and products. This shift was proposed to increase neural convergence rates, accuracy, and generalization. Echoing the development of traditional MLPs, physics-informed KAN structures were proposed shortly after, where certain knowledge of physical laws can be embedded in the training process to help the KAN converge to a physically meaningful solution [24, 25, 26]. The inference benefits of KANs were additionally demonstrated to extend to dynamical system modeling in the Kolmorogorov-Arnold Network Ordinary Differential Equations (KAN-ODEs) framework [27], where KANs replaced MLPs in the neural ODE algorithm [19], and have since been demonstrated in a variety of settings, including predator-prey dynamics, shock formation, complex equations, phase separation [27], personalized cancer treatment [28], and flashover prediction [29]. In short, a KAN network serves as the gradient getter, while a standard ODE solver integrates the solution profiles as in a traditional numerical solver. KAN-ODEs were shown to retain all major KAN benefits while also accessing the dynamical system inference capabilities of the neural ODE framework, which would appear to lend them to efficient chemical kinetic system modeling. However, a few key questions remain. KAN-ODEs have so far only been tested in relatively small systems (up to two-dimensional state variables), making their applicability and performance in larger, practical combustion systems unknown. Additionally, KANs in general have been shown to suffer substantially when trained with noisy datasets [30]. While we theorize that their direct coupling to ODE integrators combined with their sparse parameterization and smooth activations should provide KAN-ODEs with strong robustness to noise regardless of previously reported issues in generic KANs, this has not yet been studied quantitatively.
In this work, we aim to develop a chemistry KAN-ODE (ChemKAN) framework for chemical kinetic modeling by designing the Kolmorogorov-Arnold network gradient getter to learn the relationship between the current thermochemical state and the chemical source terms. Developed across two case studies of increasing complexity, the ChemKAN framework contains a physics-informed, two-stage training process that enforces the direct coupling between species production and heat release, and additionally contains a soft constraint for species element conservation. We further stabilize the optimization problems by implementing forward sensitivity analysis. Stiff chemistry is fully resolved via an attached numerical ODE solver. We study two cases here to explicitly probe the key behaviors and gaps in the model inference and solver acceleration literature identified above, both of which are addressed with ChemKANs. First, we demonstrate the capability of ChemKANs to extract realistic and multi-species models from synthesized experimental data in a comparison against DeepONets (DONs) in biodiesel production modeling. In this case, increasing levels of noise in the training data test the ability of the two different approaches to extract the true underlying behavior, and evaluate the robustness of KAN-ODEs to noisy data in light of the recent work suggesting the weakness of KANs in the presence of noise [30]. Second, we demonstrate ChemKANs as an efficient and time-saving surrogate model in an even larger system by learning zero-dimensional hydrogen combustion behavior using homogeneous reactor data that is subject to stiff dynamics, in a study designed to facilitate direct comparison against the MLP-based ChemNODE structure [17] that this second ChemKAN application is modeled after. In contrast to ChemNODE, where a reduced subset of the thermochemical state (excluding H, HO2, and H2O2) was learned using separately trained, non-interacting networks, we learn all species and temperature profiles here with a single compact ChemKAN network while retaining similar computational acceleration and performance. A successful demonstration of these tasks represents both an advancement in efficient and expressive machine learning modeling techniques for combustion, and an extension of KAN-ODEs as ChemKANs in systems significantly larger and more challenging than any previous KAN-ODE targets.
2 Methods
This section describes the mathematical details in the newly developed chemistry Kolmogorov-Arnold network ordinary differential equations (ChemKANs). We begin with a review of a chemical kinetic model in Sec. 2.1, followed by existing MLP-based kinetic modeling techniques in Sec. 2.2. Then, we discuss the implementation of the ChemKAN framework for multi-purpose chemical model inference and computational acceleration in Sec. 2.3. That subsection includes a physics-enforced ChemKAN architecture, optimal learning strategies, and physics-informed loss functions. Finally, we provide the kinetic models of biodiesel pyrolysis and hydrogen-air combustion for demonstration of ChemKANs in Sec. 2.4.
2.1 Chemical kinetics model
Given the state variables where is the temperature, is the mass fraction, and is the number of species, the net production/consumption rate for each species can be expressed as
where is the time, is the density, is the molecular weight, and is the molar production or consumption rate [31]. While the specific functional form may differ across systems, is typically a strong function of temperature (for example, in the Arrhenius form ). In some cases, Eq. 1 sufficiently describes the chemical process where heat release or consumption is negligible. However, in other cases such as combustion and pyrolysis processes, chemical reactions entail exothermic and endothermic behaviors. In these cases, the temperature of a system can be tracked with energy conservation as in Eq. 2.
where is the mixture-averaged specific heat and is the enthalpy of species . Here, the strongly nonlinear coupling of Y and T in Eqs. 1 and 2 often leads to modeling challenges such as numerical stiffness. We can express Eqs. 1-2 as a generic system of equations such that
Therefore, the thermochemical states can be predicted by integrating with an ODE integrator over time.
2.2 MLP-based models
Recent machine learning approaches for model inference and computational acceleration often rely on neural networks constructed from Multi-layer perceptrons (MLPs). Among them, we introduce two mainstream models based on deep operator networks and neural ordinary differential equations. We will use these two MLP-based models to highlight the performance of our ChemKANs (to be introduced in Sec. 2.3).
2.2.1 Deep Operator Network (DeepONet)
As one of the more popular architectures for combustion applications [21, 14, 22, 18], DeepONets [20] learn the chemical kinetic system (outlined in Sec. 2.1) through a physics-inspired separation between the system’s parameterization and the solution coordinate (Fig. 1(A)). Specifically, the solution is learned using two neural networks (branch net for thermochemical states and trunk net for time ), as per
More specifically, given an initial condition , the DeepONet reports the solution at a given time as the element-wise product () of the branch network MLP evaluated on the initial condition and the trunk network MLP evaluated on the current time, with an optional final MLP layer for additional nonlinear encoding. are the learnable parameters for the respective neural networks. The DeepONet here, as well as in other augmented structures from higher-complexity implementations such as in [22] or [18]), directly learns the solution state at future times. By eliminating the integration step, these methods have shown substantial acceleration of computational times.
2.2.2 ChemNODEs
In contrast to the DeepONet approach, ChemNODE [17] aims to learn the source terms in Eq. 3, rather than the direct thermochemical states over time . However, training challenges arose [17] when testing a single MLP network with inputs and outputs ( for and 1 for ). This led to development of a segregated model for each thermochemical state , as per
Therefore, ChemNODE constructs MLP networks, each of which reads the entire current thermochemical state vector . The network has its own unique set of parameters , and is responsible for computing the current temporal gradient of that scalar . Notably, ChemNODE is still trained via MSE loss computed against the integrated solution profile (rather than the gradients that the MLPs directly output), thanks to the differentiatble Neural ODE framework [19]. This latent-dynamics modeling for allows ChemNODEs to compute rate terms without restriction to a specific instance. This contrasts with DeepONets which require an initial condition fed into the branch network. Therefore, ChemNODEs can be effortlessly generalized as the interpretable source terms even for large-scale simulations.
We remark again that the split networks in this framework appear inefficient. In fact, this multi-network approach requires an -step training process, where each network is trained with the remaining thermochemical scalars held frozen. While this separated training strategy facilitates model convergence (and in fact was found necessary in Ref. [17] to converge the MLP gradient getters), it increases training cost exponentially with the number of species, especially when considering that no knowledge is shared between networks, even for common reactants. See Fig. 1 for visualizations of these two existing methods, as well as a simplified depiction of the method we will propose in the next sections.
2.3 Chemistry Kolmogorov-Arnold network ordinary differential equations (ChemKANs)
2.3.1 Vanilla KAN-ODEs
The KAN-ODEs framework was proposed by Koenig et al. [27] to model a dynamical system in the form of differential equations, where the gradient function is replaced by a KAN network of layers,
where KAN is the KAN representation of the system equation, parameterized by . The original KAN structure [23], so-called AddKAN, connects an -sized input to an -sized output with the learnable activation function matrix such that
where each is a unique learnable activation function connecting a single input to a single output (thus ). In other words, each input is connected to each output with a unique learnable activation function (much like in an MLP, where each input is connected to each output with a unique learnable weight), leading to a total of activation functions connecting the and layers. We use RBF basis functions in the current work as in [32, 6, 33], although the choice of is is flexible, and many other options have been proposed in the literature including B-splines [23], ReLU functions [34], and various other combinations [35, 36]. The AddKAN structure has an inherent problem of expressing the equation only with additive operations, limiting its concise expressivity for multiplicative operations. Therefore, recent studies have proposed new layer structures to address this issue and improve parameter efficiency [37, 33]. Here we introduce LeanKAN [33] which has shown promise to be the most effective and efficient for both additive and multiplicative operations. The LeanKAN structure is achieved by summation of two separate terms, and , such that
where is the multiplication hyperparameter that dictates the number of multiplication input nodes, and are the same univariate activation functions used in AddKAN (to be defined below). Overall, the LeanKAN formulation takes the matrix of activated inputs, computes sums and products in parallel according to to construct the intermediate , and finally sums the multiplication and addition components of to arrive at the next layer. Further derivation and implementation details are available for KANs in [23], KAN-ODEs in [27], and LeanKAN layers in [33].
Finally, the activation functions themselves can be expressed with gridded basis functions [32] as per
where we define as the grid size, or number of superimposed basis functions used to construct a single activation. and are the learnable network parameters that make up , where scales each gridpoint’s RBF basis functions within the sum, and scales a single base activation function . and are the index of the layer ’s matrix in Eq. 7 to which the weights are applied, while the grid itself is defined by its individual gridpoints and gridpoint spacing (or RBF spreading parameter) . While the original work of [23] suggests periodically re-gridding these basis functions to ensure they are learning on the proper input range, here we instead normalize at each layer input as in more recent works [38, 39, 27, 33] for computational efficiency. The single base activation term in each is a Swish activation function [40].
2.3.2 Novel chemistry KAN (ChemKAN) architecture
Here, we design a novel ChemKAN architecture through a unique composition of AddKAN and LeanKAN layers that shows invariance to the number of species by combining all model behavior into a single network architecture (see Fig. 1). The standard KAN-ODE architecture has the same dimensions of inputs and outputs ( thermochemical states in our current problem). In contrast, the ChemKAN architecture mimics the structure of the actual governing equations in Eqs. 12 to redirect the kinetic and thermal inputs and outputs in a physics-inspired manner. We define the full and species-only state variables with , and separate the kinetic and energy equations such that
We define the terms in these equations line by line throughout this paragraph. In Eq. 13, we encode prior knowledge of the true species production/consumption rate relationship of Eq. 1 into the ChemKAN by computing the species-only production rate from the entire state input , via a KAN network (KAN) parameterized by . Then, from the true energy equation (Eq. 2), we recognize that the temperature rate is a simple linear combination of the species production rates, with scaling factors defined by the enthalpy and specific heat values (or alternatively, in Eq. 14 by the scalars in the linear mapping parameterized by ). Following from this, the crux of the thermodynamic superstructure is a computationally trivial, simple linear sum of the already-evaluated outputs of the kinetic core, as shown in Eqs. 14 and 15. One level deeper, we recognize that a secondary effect in the true Eq. 2 is the dependence of the thermophysical parameters, specifically , on the temperature and species mixture. The error stemming from the first-order Linear() approximation’s failure to account for such thermophysical parameter variation is reflected in the term of Eq. 14, which in the final formulation of Eq. 15 is accounted for via a supplemental single-layer, single-output KAN correction carrying forward a functional dependence on the species and temperature inputs, parameterized by . Overall, ChemKAN is composed of a kinetic core structure and a thermodynamic superstructure that strongly mimic the true governing equations, operate largely in series, and include full sharing of all reaction and species production information. This architecture allows for versatile and flexible modeling by turning “on” or “off” the energy equation for standalone, kinetic core-only modeling or combined kinetic and thermodynamic modeling.
2.3.3 Network structure details
The ChemKAN structure proposed here is combined with the KAN-ODE framework detailed in Sec. 2.3.1 to leverage both the physics-guided ChemKAN architecture and the proven dynamical system benefits offered by the KAN-ODE integration. For nomenclature consistency, we henceforth refer to both the newly designed network structure as well as the complete integration with an ODE solver as ChemKAN. For the relatively larger kinetic core responsible for learning all reaction information, we define the internal layer structure as
a two-layer structure as per the notation introduced in Eq. 6. The first is an AddKAN layer [23] (which is equivalently the LeanKAN with [33]). The second is a LeanKAN layer with to inject the multiplication operator. Further discussion of this specific two-layer form and its merits are provided in detail in Ref. [33]. The thermodynamic correction structure simply comprises a sum of univariate activations evaluated on the components of ,
As in ChemNODE, differentiable ODE solvers are leveraged to enable ChemKAN encodings of the state gradients to be trained on the integrated state profiles.
2.3.4 Loss function
We define the loss function used for ChemKAN training as
Here denotes thermochemical states normalized to the [0, 1] window by subtracting the minimum then dividing by the range, while and , respectively, are the network prediction and training data. In the first MSE term, when training the kinetic core, as only species profiles are learned. For the thermodynamic superstructure, is used to train the added temperature output. We also provide an optional element conservation physics-informed loss term, or PINN term [41], similar to Ref. [42] to encourage the ChemKAN to find models that obey physical laws. There, is the number of elements in the data (i.e., H, O, N), is the number of datapoints in the temporal profiles and is the number of species represented in the data. For a given element , the element conservation term computes the mass fraction difference across all species between the training data and ChemKAN reconstruction (). This is then converted to an elemental difference via , the atom count of element in species ; , the atomic mass of element ; and , the molar mass of species . This is computed across all elements , and finally weighted by (here ) before summing with the standard MSE term. In the examples below, we use the MSE loss only for all biodiesel model inference results, and MSE loss with the PINN term for H2 model acceleration.
2.3.5 Training Process
We highlight again that this architecture incorporates significant inductive bias by separating the inference of the kinetic behavior and thermodynamic behavior into distinct layers and activations within the complete structure, giving the network an explicit functional coupling between the temperature rate and the species rates to ease its training burden while retaining an accurate model of the real exothermic behavior of the system. In itself, this serially-run inductive split should theoretically ease some of the nonlinear training difficulties observed in [17] while retaining a single cohesive and information-sharing structure. Additionally, and more importantly to stable training and convergence behavior, it allows us to separate the training itself into a kinetic and a thermodynamic stage to greatly reduce the inference burden, as discussed here:
1.
Training stage 1—core kinetics : All inputs are used, and only the species production rate outputs are computed (see Eq. 13). The thermodynamic superstructure of Eqs. 14 and 15 is not used in this stage, and the input temperatures are simply read in from the training data to provide the kinetic core of the network with a simpler training task. See the grey highlight in Fig. 2. For cases without heat release, this step in isolation is sufficient for a complete model. With heat release, we move to stage 2 once stage 1 is converged.
2.
Training stage 2—thermodynamic superstructure and : Once stage 1 is converged, the thermodynamic superstructure is added and the temperature rate is explicitly learned. The entire network (Eq. 13 stacked with Eq. 15) is updated in order to infer the temperature together with all species. See the red highlight in Fig. 2, which is stabilized during training with the already-converged behavior of the grey kinetic core.
This two-stage training process contrasts with the stage training process of ChemNODE [17], which requires networks to all learn their own distinct representations of what we know to be shared kinetic and thermodynamic governing laws. The current ChemKAN approach, while trained in two distinct stages, has an overall structure resembling a simple feedforward KAN thanks to its stacked design where all kinetic and thermodynamic information is shared across all inputs and outputs. For downstream evaluation, a single forward pass through the combined network of Fig. 2, or alternatively through Eqs. 1315, predicts the complete thermochemical state vector .
The learnable parameters are trained by the Adam optimizer [43]. While adjoint sensitivity analysis was used in the original KAN-ODE paper [27] to compute the gradients of a loss function , in the current work we note the numerical instability present in many chemical kinetic modeling problems due to numerical stiffness [44, 22, 18, 17, 3], potentially leading to failures with adjoint sensitivity analysis in the training process [3]. To prevent this, we implement forward sensitivity analysis to mitigate this potential stiffness issue in the ODE solver [3]. The learning rate was set to .
2.4 Data generation for case studies
We introduce two chemical reaction systems to demonstrate ChemKANs as a multi-purpose modeling technique. The first case of biodiesel production will illustrate the effectiveness of ChemKANs for inferring a chemical kinetic model purely from data. Then, we show computational acceleration for high-fidelity simulations using ChemKANs with an example of hydrogen-air combustion subject to significant numerical stiffness.
2.4.1 Model inference from noisy data – biodiesel production
Biodiesel production involves the transesterification of branched triglyceride molecules (TG) with methanol into straight-chain methyl ester molecules, described in Ref. [45] and previously modeled in Refs. [5, 46]. A motivating goal here is to evaluate the ChemKAN framework’s capability to learn the true models underlying experimental data in a preliminary and numerically well-behaved case related to combustion chemistry, and to probe its robustness to noisy data masking the true underlying signal.
With the three byproducts di-glyceride (DG), mono-glyceride (MG), and glycerol (GL), the three-reaction system can be expressed as
where the reaction rates scale with temperature via the standard Arrhenius law as , with for the three reactions. As in Ref. [5], we generate data in this case using kcal/mol and , with isothermal experiments at temperatures randomly sampled in the range of K to K. We define the species scalar quantities here as concentrations rather than mass fractions, to match convention with the governing equations. Initial TG and ROH concentrations are randomly sampled uniformly between and with all other intermediate and output species initialized at zero. 20 training data sets and 10 testing data sets are generated, with 30-second time windows in both consisting of sampled points. The temperature-dependent yet isothermal reaction rates present in this system motivate use of the kinetic ChemKAN core structure only. In other words, we drop the thermodynamic superstructure of Eq. 15.
In this case we additionally probe the effectiveness of ChemKANs in the presence of significant experimental noise. To do so, we add increasing amounts of noise to the data and evaluate the ChemKAN’s performance against that of a standard Deep Operator Network [20].
2.4.2 Model acceleration – hydrogen-air combustion
Despite its relatively simple chemistry, Hydrogen-air combustion is subject to stiffness, leading to high computational costs and numerical instability in the ODE solver. The main goal here is to achieve high accuracy in species and temperature profile reconstruction while reducing computational cost compared to the detailed chemistry solver using ChemKANs. We generate training data in Cantera [47], using the H2/O2 mechanism from GRI-Mech 3.0 [48] (9 species, 29 reactions). Data are generated at all 36 combinations of initial temperatures in {950, 1000, 1050, 1100, 1150, 1200} K and equivalence ratios in {0.7, 0.9, 1.1, 1.3, 1.5}, with the [1150 K, 1.3] case withheld as unseen testing data. The problem setup here is largely identical to that studied originally in ChemNODE [17] to facilitate fair comparison, although here we add the single withheld testing dataset to probe the ChemKAN’s robustness (thus 35 training datasets are used, compared to the 36 in ChemNODE).
In addition to significantly increased numerical stiffness and complexity, a key fundamental difference between the current combustion system and the previous biodiesel synthesis case is the presence of substantial two-way temperature coupling as per Eqs. 12. To account for this, we include the thermodynamic ChemKAN superstructure and the two-stage training process outlined in Secs. 2.3.22.3.5.
3 Results and Discussion
3.1 Biodiesel model inference from noisy data
We begin with analysis of the biodiesel model inference of the ChemKAN (the kinetic core only for this isothermal case), and a comparison against a traditional DeepONet for an identical task. DeepONets were selected as the target of comparison due to their general use in the scientific machine learning community, to provide a baseline against which to evaluate our first ChemKAN results. Note that comparison against ChemNODE is saved for the hydrogen-air combustion case in Sec. 3.2.
3.1.1 Model performance of the proposed ChemKAN architecture
The kinetic core (Eq. 13) of the ChemKAN comprises two layers: one AddKAN and one LeanKAN, as discussed in Sec. 2.3.2. Note that the thermodynamic superstructure (Eq. 15) is not applied in this problem as the process is isothermal. A four-node hidden layer is used, with all activations comprised of three-point grids. The multiplication hyperparameter in the LeanKAN layer is set to 2 as per the recommendation of Ref. [33]. The training and testing data are sampled from the ground truth at a sparse time interval of 1 s, leading to 30 total data points for each individual species. The ChemKAN is trained on this data for epochs. Figure 3 shows example time-history species profiles for an (unseen) test case with the ground truth, noisy test data, and prediction by the learned ChemKAN model. With clean, noise-free data, the ChemKAN successfully learns the underlying model, showing good generalizability with accurate predictions at the training time steps, and smooth profiles in between.
A common trait of experimental data from which machine learning techniques are tasked with inferring models is experimental uncertainty or noise, which is another avenue in which even properly parameterized deep learning techniques may overfit as they struggle to distinguish between genuine underlying trends and experimental noise. In fact, this is a known problem with standard KAN networks, where it was shown in Ref. [49] that one of the original cases in Liu’s original KAN work [23] fails with a small amount of added noise, a phenomenon that was further studied in more quantitative detail in Ref. [30]. In the current work, we further probe whether coupling to inherently noise-robust ODE solvers helps ChemKANs to extract useful models from increasingly noisy data. To do so, we task ChemKANs with extracting models from the same dataset with varying amounts of noise added (up to 15%) as shown in Fig. 3. Surprisingly, the ChemKAN, even with significant amounts of noise, demonstrates strong robustness and a capability to infer smooth and accurate solution profiles that correspond to the underlying true data. We provide detailed discussions on how ChemKAN performs well with comparison against DeepONet in the following subsections.
3.1.2 Neural scaling with noise-free data
Neural scaling is an effective method of measuring parameter efficiencies in neural networks. Here different parameter sizes were investigated by changing the number of nodes in the hidden layers. Training and test losses were evaluated for data without noise to isolate the underlying expressive capabilities of each technique. Figure 4 reveals neural scaling at orders of 1.0 and 0.6 for the training and testing metrics of ChemKANs (where the order of neural scaling is defined as the power to which the loss decreases with respect to the number of parameters). While we might expect values up to 4 based on prior KAN [23] and KAN-ODE [27] studies, we remark here that the low-order neural scaling appears to indicate relatively saturated training of the ChemKAN with errors already around with just 72 parameters, rather than poor convergence which might otherwise have been indicated if low convergence rates were coupled with poor loss metrics.
To further probe the nuances of the ChemKAN’s convergence efficiency, we compare its results with DeepONet. The ChemKAN iterates roughly an order of magnitude slower but converges in fewer epochs (as originally discussed in [27]). Thus, to facilitate fair comparison the ChemKAN was trained only for 5,000 epochs, while the DeepONet was trained for 50,000 epochs (10 times more than ChemKAN). Two key distinctions between ChemKAN and DeepONet are as follows. First, as discussed earlier, the extremely sparse ChemKANs (toward the left half of Fig. 4(A)) are able to reach remarkably low error even with just 78 parameters, while the DeepONets see significantly worse performance at sparse parameterizations (i.e., the left half of Fig. 4(A)). Secondly, we note that the DeepONet sees significantly higher order neural convergence in the training results, allowing it to surpass the training performance of the ChemKAN at above 200 parameters, with remarkably strong training accuracy for the largest, 456-parameter DeepONet studied here. Linear fits with slopes are shown in Fig. 4(A) to illustrate this point, where the last ChemKAN and last two DeepONet points are excluded as they begin to plateau in training loss. From this observation, a large-enough DeepONet seems to outperform ChemKAN when looking only at the training losses.
To further contextualize the structural efficiency, we plot results for the testing error of the same neural convergence runs in Fig. 4(B). Here, we notice a significant departure from the training convergence rates in Fig. 4(A). Unlike the neural convergence for training data, the ChemKAN is seen to outperform the DeepONet at all sizes, with the largest ChemKAN leveling out and retaining nearly the same testing performance as the second-largest ChemKAN. The DeepONets, while enjoying a faster neural convergence rate below 308 parameters, notably fail to plateau and instead appear to diverge at higher parameter counts. When compared against the training results in Fig. 4(A), we observe two distinct modes of training saturation. Saturation, or the point where the linear fit no longer holds, appears to occur for the two largest DeepONets. For the ChemKAN, we might either interpret the entire profile to be saturated, or highlight the single largest network as the saturation point. Regardless, what we observe in these high-parameter networks is high robustness in the ChemKAN to overfitting (with a flat testing loss plateau), compared to the significant overfitting and divergence seen in the DeepONet past 308 parameters (i.e., the last two points).
It is unsurprising that a standard deep learning technique begins to overfit a small dataset when given a large number of parameters. The DeepONet overfitting past saturation leads to further decreases in training loss accompanied with significant increases in testing loss. What we do find surprising, however, is the ChemKAN’s apparent resilience to overfitting, even with similar parameter counts. In this context, its original failure to reach the same training performance as the DeepONet appears to be a strong point of the method rather than a drawback, as it reaches a minimum value for both training and testing and then remains robust to superfluously added parameters, while the DeepONet clearly requires additional care to avoid overfitting. This pilot neural convergence test suggests that ChemKANs are robust to overfitting, and motivates further study of their capability in a second, more realistic model inference scenario.
3.1.3 Benefits of latent-dynamics modeling: noisy inference comparison
To gain deeper understanding of how ChemKAN outperforms DeepONet, we further test training behaviors of ChemKAN and DeepONet with noisy data. Early stopping is carried out for both network structures at 10,000 epochs, to enable fair comparison and limit overfitting in the DeepONet cases. In all cases, a DeepONet with 308 parameters was used to compare against a ChemKAN with 156 parameters. The 308-parameter DeepONet was chosen based on the results of Fig. 4, where this was seen to be the largest DeepONet before the test loss began overfitting. The 156-parameter ChemKAN, meanwhile, was chosen to roughly match the DeepONet’s training and testing performance at zero noise. In other words, we chose the best-performing DeepONet size possible based on the preliminary neural convergence study, and then sized the ChemKAN according to the zero-noise performance of both networks. In more detail, the DeepONet had a three-layer branch network with eight nodes per layer and a two-layer trunk network with seven nodes in the first layer and eight in the second layer, with a final output layer converting these two eight-dimensional layers to the six-dimensional solution vector. The ChemKAN, meanwhile, had a single hidden layer with four nodes, two of which included multiplication operators () as per the recommendation of [33], and three gridpoints per activation.
Figure 5(A) shows average training results after epochs across the 20 training and 10 testing cases at varied noise levels from to . As discussed in Sec. 3.1.2, the training MSE with 0% noise shows that the 308-parameter DeepONet slightly beats the 156-parameter ChemKAN in training performance, and is slightly worse in terms of reconstructing the unseen testing data.
As increasing noise is added to the system, we see in the standard training and testing metrics of Fig. 5(A) that both networks unsurprisingly see increases in training and testing errors. Roughly, the increase in MSE error in the training losses scales with the square of the noise, as follows from Eq. 19. For example, the ChemKAN sees an increase in MSE between 0% noise and 1% noise of , where the second degree scaling suggests that a 25 larger increase of might be expected between 0% noise and 5% noise. This is indeed observed, with an increase in this latter case of . In other words, the increase in training error for both networks as noise is added can be attributed to the effect of the noise itself on the loss function (Eq. 19), and does not appear to indicate any problems with the two networks’ capabilities to fit the increasingly noisy training data. For a direct comparison, The ChemKAN retains lower testing error throughout all tested noise values, and at noise and above is actually able to reach a lower training error than the DeepOnet. Looking at the big picture, however, results in these two metrics remain within a factor of two of each other at all noise levels, indicating largely similar performance.
Upon further evaluation, we found that the training and testing MSEs evaluated on noisy datasets do not fully capture the effects of noise on useful model predictions, as overfitting can occur not only to the training conditions but also to the noise present in the data. To more effectively compare these frameworks, we introduce a noise-free MSE metric (similar to Ref. [30]),
which differs from the test MSE in Eq. 19. This noise-free metric serves to quantify the capability of the two modeling approaches to accurately extract the true underlying model from noisy data, rather than overfit the noise or otherwise fail to deliver a useful model. With this metric, much more significant performance shifts can be seen in the noise-free testing values in Fig. 5(A), which are the true indicators of the two approaches’ capabilities to extract hidden underlying signals from noisy data. While the added noise clearly increases the training and testing error when evaluated on the noisy data itself for both model types, the impact of the added noise on the noise-free testing error for the ChemKAN is relatively small (a roughly increase from to noise), suggesting that it is able to extract the true underlying behavior well from the noisy datasets. We reiterate the significance of this result in the context of Ref. [30], where standard KANs were demonstrated to fail when faced with small added noise (while here the ChemKAN retains good performance). In contrast, the DeepONet sees a 5 increase in noise-free testing MSE from to noise, with its final noise-free testing loss 4.4 larger than that of the ChemKAN.
The loss profiles in Figs. 5(B) provide further insight on the training dynamics that lead to this significant discrepancy in noise-free testing reconstructions. In the noise training cycle of Fig. 5(B), we see fairly standard behavior across the training and noise-free testing traces for the DeepONet and ChemKAN, with all quantities steadily decreasing for the entire duration of training. This is the expected result, as we have sized the DeepONet based on Fig. 4 to avoid any overfitting in the noise-free case. As we increase the amount of noise to , the training loss values are heavily penalized (due to the noisy data used in the computation of Eq. 19), while the noise-free testing values are slightly penalized but remain comparatively strong, indicating that both approaches are at least to a certain extent able to extract the true underlying model from the noisy data. While the training loss drops quickly and then plateaus in all four subplots, we see in the 7% noise case of Fig. 5(B) that the DeepONet noise-free testing loss dynamics begin to suffer, with a minimum value near 5,000 epochs and a slight upward trend toward later epochs, likely due to overfitting. In the 15% noise case of Fig. 5(B) this issue is further exacerbated, with an early minimum near 1,000 epochs followed by significant overfitting to the noisy data for the remainder of the training profile. The ChemKAN in both cases remarkably continues to drop its noise-free testing loss even while the DeepONet is overfitting, with late-epoch dynamics showing plateaued minimum values rather than the overfitting seen in the DeepONet. This echos the behavior seen in Fig. 4, where the ChemKAN did not overfit and instead simply plateaued at its minimum training and testing errors.
Reconstructed training data profiles are shown for the 15% noise case in Fig. 6. While the DeepONet and ChemKAN are both roughly able to find the unseen, noise-masked ground truth profile, a close inspection reveals not only better ChemKAN fits, but also notably jagged profiles from the DeepONet as it attempts to overfit the noise present in the training data. These results suggest that the dynamical system exploitation inherent to the ODE-based framework of ChemKANs (and KAN-ODEs in general) are able to mitigate or even entirely resolve the issues observed in Ref. [30] regarding noisy data with KAN networks.
In summary, we have demonstrated ChemKANs as a promising tool for model discovery in temperature-dependent chemical kinetic systems, especially with realistically noisy datasets. They show promise not only on the scientific side of the problem, where they were demonstrated here to have significant capability compared to a standard tool in discovering models hidden under synthetic noise, but also on the machine learning side of the problem, where we have demonstrated that the neural ODE implementation of KAN-ODEs and ChemKANs helps them to overcome the noisy data limitation recently shown in vanilla KAN networks [30].
3.2 Hydrogen combustion acceleration from homogeneous reactor data
In our second case study, we investigate the use of ChemKANs as a reduced-order solver acceleration framework for known chemical models. We hypothesize based on previous high-order neural scaling results [27, 23] as well as the strong performance at low parameter counts seen in Fig. 4 that the use of a KAN structure instead of an MLP will allow for similar accuracy in network predictions with fewer parameters and lower cost. We additionally aim to evaluate the novel network architecture of the ChemKAN, and whether its physics-based, single-network structure is able to model behavior that required MLPs in the Neural ODE framework of Ref. [17]. To reduce computational cost to the furthest extent possible, we use a single hidden layer of just three nodes for the core kinetic network (Eq. 13). This three-node hidden layer works as a latent representation of the system’s dynamics, compressing the information from 10 thermochemical states in the input and 29 reactions in the kinetic model. Leveraging our knowledge of the functional form of the governing equations (Eqs. 1 and 2) and their strong multiplicative behavior, we use here, defining all three hidden nodes using the multiplication operator.
We begin in Fig. 7 with a demonstration of the ChemKAN homogeneous reactor reconstructions for one training case (top row, and K) and the unseen testing case (bottom row, and K), after the two-stage training process outlined in Sec. 2.4.2. These reconstructions were generated entirely by the ChemKAN given only the initial conditions. Overall the learned model successfully predicted the temperature and mass fractions in both cases, with no notable deterioration in the testing case (as expected from the strong testing results and robustness to overfitting observed in the previous biodiesel investigation). We further emphasize that the ChemKAN was largely able to capture the behavior of the low-concentration and highly reactive species , , and (Fig. 7(C1-C2)) that were neglected in ChemNODE [17].
A broader comparison is shown in Fig. 8(A), where the MSE (in the normalized u units) is plotted across not only the initial set of 35 training and 1 testing initial conditions, but a wider set of 441 total initial conditions (406 of which were unseen during training) at a finer resolution in the same range. The low-temperature initial conditions see near-perfect reconstructions, as the ignition delays there are larger than the studied time window, leading to smooth gradients and easily trainable, near-isothermal behavior in those cases. In the remainder of the domain, strong performance is seen at all training conditions, and additionally at the single testing condition plotted in Fig. 7. In terms of generalization to intermediate temperatures and equivalence ratios, the ChemKAN performed very well throughout the vast majority of the domain, with many testing points even surpassing the accuracy of nearby training points at and above 1050K. We notice that toward the slower-igniting cases, however, the ChemKAN struggled more with generalization. At 1000K, for example, all six training points saw strong MSE values in the range, although a few of the intermediate equivalence ratios suffered. At 987.5K, however (one tick below 1000K), all testing points saw poorer performance in the range. We can conclude from these results that the ChemKAN retains a strong capability to generalize in the more challenging H2-air combustion case (as was originally reported in the biodiesel modeling case), but with practical limits. At higher temperatures it was remarkably even seen to outperform training results in broad swaths of unseen testing conditions, but in the more initiation-sensitive low-temperature cases, it was only able to perform well locally near the training points, indicating that denser training data is needed to optimize performance in those regions.
We finally plot the actual and KAN-ODE predicted ignition delays in Fig. 8(B), for the 30 studied cases that ignited (the lowest temperature cases did not see ignition given the time span of 0.6 s). Ignition here is defined as the point of maximum temperature rise rate, as in Ref. [17]. Accuracy is strong across the board, even in the testing case.
This collection of results shows that the ChemKAN structure was able to accurately learn the dynamics of all nine species and temperature scalars across the same set of initial reactor conditions as was studied using traditional MLP-based Neural ODEs in ChemNODE [17]. Compared to the six species plus temperature scalars learned there via seven unique MLP networks with 91 parameters each according to standard MLP parameterizations (637 total parameters), the current ChemKAN was able to learn the complete set of thermochemical scalars (nine species plus temperature) using a single, 344-parameter network. While training took place in two stages to decouple the kinetic and thermodynamic behavior and facilitate convergence, the final network remains a single cohesive structure with shared information across all nodes, eliminating the redundancies in repeated yet isolated 91-parameter MLPs.
Finally, regarding computational efficiency, we report that the average time to solve all 36 homogeneous reactor conditions 100 times each (3600 total solves) in the Arrhenius.jl combustion solver package [50] was 2 faster on average when switching from the detailed chemistry to the reduced ChemKAN framework. By learning the relationship between the current thermochemical state and the chemical source terms, the ChemKAN is capable not only of predicting ignition delay times and homogeneous reactor solution profiles 2 faster than the detailed model, but also of generalizing to other simulation conditions including simple laminar flames and complex 2-D and 3-D turbulent combustion conditions. Such downstream uses of similar surrogate machine learning models were discussed and tested in Refs. [16] and [17], where a 2 speedup in the chemical solver (which is often the most computationally expensive component in a reacting flow simulation) implies the potential for substantial acceleration unlocked by ChemKANs while retaining the full-sized solution state vector. A summarized comparison of ChemKAN and ChemNODE is provided in Table 1.
| ChemNODE | ChemKAN | |
| No. Networks | 7 | 1 |
| Parameters | 637 | 344 |
| Species modeled | H2, O2, H2O, N2, O, OH | H2, O2, H2O, N2, O, OH, H, HO2, H2O2 |
| Speed-up vs. full chemistry | 2.3x | 2.0x |
3.3 Current limitations and areas for further research
While we have demonstrated that our ChemKANs provide remarkable inference capability and expressivity, there remain drawbacks. The total speedup reported in Table 1 is ultimately not as large as we believe that it could be. While still more than competitive with that of ChemNODE especially considering the complete thermochemical source term with low-concentration radicals that ChemKAN provides, it appears underwhelming in light of the rest of the significant performance gains enabled by ChemKAN in other comparisons throughout this work. That being said, we believe that the reported speedup is a conservative lower bound on the potential for ChemKAN and KANs in general. The original KANs [23] are at the time of original submission of the current work less than a year old and are known in the literature to be much slower than comparable MLPs (this is even acknowledged by the authors of the original KAN paper [23]). However, extensive research is actively ongoing to resolve this issue including parameter efficiency improvements [33, 51, 35] and prediction acceleration techniques [38, 52, 53, 54, 34, 39, 55]. It’s unclear which methods will ultimately prevail, but the recent emergence of KANs and the large amount of work proposing various techniques for their acceleration suggest significant promise in the near future for substantial and relatively lower-effort acceleration, compared to the more mature and well-developed MLPs where we believe it reasonable to expect a slower pace of future development.
We have additionally in this section compared a baseline ChemKAN implementation against a baseline ChemNODE implementation. Later works exist that appear to successfully combine ChemNODE with augmented loss functions, autoencoders, and latent space time stepping. The most recent, which includes all of these techniques, is reported in Ref. [42] (“Phy-ChemNODE”). We do not draw comparisons between the current ChemKAN implementation and the larger-scale, combined-methodology results in [42], as our current aim is to compare the pure performance of the ChemKAN against the MLPs that underlie both ChemNODEs and Phy-ChemNODEs. All further augmentations carried out in Phy-ChemNODE that go beyond this baseline can be replicated with ChemKANs, and an interesting target of future studies may be on quantifying the performance gains of ChemKAN when applied in tandem with other advanced neural network structures.
4 Conclusions
This work introduced ChemKANs, a novel physics-informed machine learning technique based on the general KAN-ODE framework with a specialized structure tailored explicitly for chemical kinetic modeling and acceleration. Its two-part design allows for application in isothermal or exothermic systems through a kinetic core and an optional thermodynamic superstructure that can be trained and applied as a single cohesive framework. Model inference in a preliminary biodiesel synthesis case using only the kinetic core revealed that ChemKANs are remarkably robust to overfitting. Fair comparisons against a generic Deep Operator Network approach revealed that with increasingly bulky parameterizations, ChemKANs retained plateaud optimal loss metrics, while DeepONets saw minor further decreases in training loss accompanied with an increase in testing loss, as might be expected from a standard deep learning approach. With added noise, this difference was further exacerbated, with the DeepONet achieving low training loss through jagged fits to the noise itself, while the ChemKAN converged its training, testing, and noise-free testing performance with smooth fits to the hidden underlying data. While alone promising, these results are of particular interest given prior works in the literature that have casted doubts on the effectiveness of KAN structures on noisy functions.
In a second case, ChemKANs were demonstrated as efficient acceleration surrogates for learning chemical source terms in a hydrogen combustion case. A two-stage training process for the kinetic core and thermodynamic superstructure enabled a single, 344-parameter ChemKAN to accurately learn complete solution profiles across a range of hydrogen-air homogeneous reactor initial conditions, a significant reduction in parameter and network bulk compared to previous MLP-based neural ODE approaches that required 637 parameters to learn a truncated set of solution profiles. Timing comparisons against the detailed mechanism revealed a 2 speedup when using the ChemKAN surrogate model, which is significant for downstream applications of the hydrogen combustion surrogate learned here (for example, 3-D turbulent reacting flow). In summary, we find that ChemKANs are a promising tool for both dynamical system modeling and acceleration tasks in combustion chemistry. In doing so, we have also successfully advanced the underlying KAN-ODE framework to much larger, practical systems than have been studied previously. We hope that these promising preliminary case studies motivate future implementation of KAN layers and modules in combustion machine learning applicatins.
Acknowledgement
This work is supported by the National Science Foundation (NSF) under Grant No. CBET-2143625. BCK is partially supported by the NSF Graduate Research Fellowship under Grant No. 1745302.
CRediT authorship contribution statement
Benjamin C. Koenig: Conceptualization, Methodology, Software, Investigation, Writing - Original Draft. Suyong Kim: Conceptualization, Methodology, Writing - Original Draft, Writing - Review Editing. Sili Deng: Funding Acquisition, Resources, Writing - Review Editing.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.